1 litre mixture of CO and CO2 is taken. This is passed through a tube containing red hot charcoal. The volume now becomes 1.6 litre. The volumes are measured under the same conditions. Find the composition of mixture by volume.

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
We have
initial volume of CO+CO2 =1 lt
final volume of the mixture =1.6lt
this final volume has increased due to the more CO formed
C+CO2 ---2CO
Let the initial volume of CO2 be xlt,volume of CO=1-x lt
initial volume of CO2=x
final volume of CO=2x
therefore final volume of 1-x+2x=1+x
1+x=1.6
x=.6lt
volume of CO2=.6lt
volume of CO=.4lt
percentage of CO2=60%
percentage of CO=40%